Use two equations with two variables to solve the problem.
The sum of two numbers is 15 less than twice the first number. Their difference is 5 less than twice the second number. Find each of the numbers.
I appreciate any help that can get me to see the equation and the solution. Thank you so much
Sharon

Answer by richwmiller(17219)

sum means we are going to add things
less means we are subtracting something
is means equals
twice means two times
difference means subtraction
x+y=-15+2x
or
x+y=2x-15
Now the second sentence
x-y=-5+2x
or
x-y=2x-5

#1
x+y=2x-15
#2
x-y=2x-5
combine
y=x-15
-y=x-5
add
0=2x-20
20=2x
x=10
10-5=-y
5=-y
y=-5
(10,-5)
#1
x+y=2x-15
10-5=20-15
5=5
#2
x-y=2x-5
10-(-5)=2*(10)-5
15=20-5
15=15
ok